A group of students estimated the length of one minute without reference to a watch or clock, and the times (seconds) are lis(Reject, Fail to reject) H0. There is (sufficient , not sufficient) evidence... 60 seconds (is, is not) correct. It (does not appear, does appear) that, as a group....
Answer #1

Standard deviation, S = ECX-X)? . X=71.6 n-l 2 2(x-7)? (81-71.6)+(90-71.6997(50-71637(49-71632 +(36-71.6)+(68-71.6)+ (75-71,6P-value = P(t, df) (using p-value from = P(2.56,14) t-score calculaton 0.023 1. P-value 0.023 Here, p-value at iei P-value co

Please LIKE if it really helps you!!!